OpenStudy (anonymous):

what is the standard form of the parabola with the vertex at the origin and the focus of (0,6)

OpenStudy (anonymous):

i need a reply so i can finish my homework.

OpenStudy (amistre64):

4ay=x^2, maybe; given that a = 6

OpenStudy (amistre64):

you might have to consult your text or material for what "standard" form looks like in this context

OpenStudy (dumbcow):

standard form : \[y = ax^{2} + bx +c\] vertex (0,0) c=0 b=0 \[a = \frac{1}{4p}\] where p is distance from vertex to focus p=6 a = 1/24 \[y = \frac{1}{24}x^{2}\]